Water from the Don Juan Pond in Antarctica is 4.40 M sodium chloride. If you have a 

340. mL sample of lake water, how many grams are in solution?


1
Expert's answer
2021-03-22T06:36:03-0400

Molar Mass of water = 18.01528g/mol

Molar Mass of sodium chloride = 58 g/mol

= 4.4× 18.01528 = 79.267232

350 ml = 350 g

= (58/79.267232) ×340

= 248. 78 grams